
Here we will show you how to convert the implied probability of 2.99 percent to decimal odds. To convert the implied probability of 2.99 percent to decimal odds, we divide 1 by 2.99 and then we multiply the quotient by 100, like this:
1 ÷ 2.99 ≈ 0.334448
0.334448 × 100 ≈ 33.445
2.99% probability ≈ 33.445 decimal odds
That's all there is to it. Now you know that you need decimal odds of about 33.445 if you want a 2.99 percent implied probability of winning.
Implied Probability to Decimal Odds Converter
You can use the form below to convert another implied probability percentage to decimal odds.
Convert 3 percent implied probability to decimal odds
Here is another implied probability to decimal odds conversion that we have done for you.
